# Multi-Agent Target Assignment and Path Finding for Intelligent Warehouse: A Cooperative Multi-Agent Deep Reinforcement Learning Perspective

### v2

@@ 4.1 Modeling TAPF as a MARL Problem
Observation space: An agent’s observation contains the position and velocity itself, the relative positions of all tasks, the relative positions of the other agents, and the relative positions of neighboring obstacles. In this study, the number of other visible agents in an agent’s observation can be set to equal to or less than $N-1$. Taking Figure 1 (a) as an example, the five purple circles represent five agents, and the five black circles represent five tasks. For obstacles, we set the agent can perceive the relative positions of the adjacent obstacles, and the number of adjacent obstacles can be set to any integer. In Figure 1 (a) and (b), different color curves denote the navigation trajectories of different agents.
The physical dynamics of agents: Before describing the action space, we first describe the dynamics of agents. Note that our work is the first to consider the dynamics of agents in TAPF for the intelligent warehouse. As shown in Figure 2 (a), the ”No dynamics” schematic diagram represents previous literature [27, 33] methods that only focus on the next position of an agent, but don’t consider velocities and accelerations of agents. Further, the dynamics of agents in previous literature are solved by using traditional methods [34]. On the contrary, the ”Dynamics” schematic diagram in Figure 2 (b) represents our method that considers the dynamics of agents. We can calculate agents’ velocities and accelerations.
The output of our policy network is the magnitude and direction of four cardinal directions force ($\vec{F_{x}},\vec{F_{-x}},\vec{F_{y}},\vec{F_{-y}}$) applied to the agent. According to Newton’s second law of motion, $\vec{F_{direction}}=m*\vec{a_{direction}}$ (where $m$ is the mass of the agent), we can get the acceleration $\vec{a_{direction}}=\frac{\vec{F_{direction}}}{m}$. According to basic physics knowledge, we can get the velocity of the agent $\vec{v_{t(direction)}}=\vec{v_{0(direction)}}+\vec{a_{direction}}*\Delta t$, where $v_{0}$ is the initial velocity and $\Delta t$ denotes a time interval.
Action space: In this study, the action space of an agent is continuous, representing the movements. As shown in Figure 3, an agent obtains a velocity between 0.0m/s and 1.0m/s in each of the four cardinal directions [move_left ($\vec{v_{-x}}$), move_right ($\vec{v_{x}}$), move_down ($\vec{v_{-y}}$), move_up ($\vec{v_{y}}$)], the final action ($\vec{v}$) is calculated by the the vector sum of the four velocities.
Figure 4: Distansce matrix of tasks and agents: $\|\boldsymbol{P_{tasks}-P_{agents}}\|$. The element $dis(a\ i,t\ i)$ represents the distance between $agent\ i$ and $task\ i$.
Reward function: We define the reward function as:
@@
    wherendenotes the number of elements that satisfy
[…]
    <0.05
[…]
    0.05
[…]
  • 2.
    reward_distance_tasks_to_agents
    = -$\|\boldsymbol{P_{tasks}-P_{agents}}\|_{min}$
  • 3.
    reward_collision_agents_to_obstacles = -$2*n$,
    wherendenotes the number of elements that satisfy
[…]
    <R
@@
  • 4.
    reward_collision_agents_to_agents = -$2*n$,
    wherendenotes the number of elements that satisfy
[…]
    <2R
@@
For reward_success, $\|\boldsymbol{P_{tasks}-P_{agents}}\|$ represents the distance matrix of all tasks to all agents. Figure 4 shows a detailed distance matrix. If the distance of $agent\ i$ to $task\ i$ is less than $0.05m$, we give a $+100$ positive reward to enforce agents to navigate to the tasks. If there are $n$ elements in $\|\boldsymbol{P_{tasks}-P_{agents}}\|$ less than $0.05m$, we set reward_success = +$100*n$. Reward_distance_tasks_to_agents is created to provide a dense reward to accelerate the training speed. In $\|\boldsymbol{P_{tasks}-P_{agents}}\|_{min}$, the subscript $min$ means that we sum the minimum distance of each task to every agent.
Reward_collision_agents_to_obstacles is aim to punish agents from collision to obstacles. The definition of $\|\boldsymbol{P_{obstacles}-P_{agents}}\|$ is similar with $\|\boldsymbol{P_{tasks}-P_{agents}}\|$, $\|\boldsymbol{P_{obstacles}-P_{agents}}\|$ denotes the distance matrix of obstacles and tasks. If the distance of an agent to an obstacle is less than $R$, a negative reward $-2$ is given. If there are $n$ elements in $\|\boldsymbol{P_{obstacles}-P_{agents}}\|$ less than $R$, we set reward_collision_agents_to_obstacles = -$2*n$.
Reward_collision_agents_to_agents is aim to punish agents from collision to agents. The definition of $\|\boldsymbol{P_{agents}-P_{agents}}\|$ is similar with $\|\boldsymbol{P_{tasks}-P_{agents}}\|$, $\|\boldsymbol{P_{agents}-P_{agents}}\|$ denotes the distance matrix of agents and agents. If the distance of an agent to an agent is less than $2R$, we give a negative reward $-2$. If there are $n$ elements in $\|\boldsymbol{P_{agents}-P_{agents}}\|$ less than $2R$, we set reward_collision_agents_to_agents = -$2*n$.
4.2 Solving TAPF Problem via Cooperative MARL
In this study, we introduce MADDPG [31] to solve the TAPF problem. Considering that agents are homogeneous, thus the agents can share the same policy network. This makes learning policy more efficient. Let $N$ denote the number of agents, $\pi$ denote the policy parameterized by $\theta$. The centralized critic $Q$ is parameterized by $\phi$. The parameters $\theta$ for policy $\pi$ can be updated iteratively by the associated critic $Q$:
$\nabla_{\theta}J\left(\theta\right)$ $=\mathbb{E}_{\left(\boldsymbol{x_{t}},\boldsymbol{a_{t}}\right)\sim\mathcal{D}}\Big{[}\nabla_{\theta}\pi\left(o_{t}^{n}\mid\theta\right)$ (3)
$\quad\nabla_{a_{t}}Q\left(\boldsymbol{x_{t}},a_{t}^{1},\ldots,a_{t}^{N}\mid\phi\right)|_{a_{t}^{n}=\pi\left(o_{t}^{n}\mid\theta\right)}\Big{]}$
where $\boldsymbol{x}_{t}$ and $\boldsymbol{a}_{t}$ are the concatenation of all agents’ observation and action at time step $t$, $\boldsymbol{x}_{t}=\left(o_{t}^{1},\ldots,o_{t}^{N}\right)$ and $\boldsymbol{a}_{t}=\left(a_{t}^{1},\ldots,a_{t}^{N}\right)$. $o_{t}^{n}$ is the observation received by agent $n$ at time step $t$. $\mathcal{D}$ denotes the replay buffer containing tuples ($\boldsymbol{x}_{t},\boldsymbol{a}_{t},r_{t}\cdots$).
@@ 4.2 Solving TAPF Problem via Cooperative MARL
$L\left(\phi\right)=\mathbb{E}_{\left(\boldsymbol{x}_{t},\boldsymbol{a}_{t},{r}_{t},\boldsymbol{x}_{t+1}\right)\sim\mathcal{D}}\left[\left(Q\left(\boldsymbol{x}_{t},\boldsymbol{a}_{t}\mid\phi\right)-y_{t}\right)^{2}\right]$ (4)
[…]
$y_{t}=r_{t}+\gamma Q^{-}\left(\boldsymbol{x}_{t+1},a_{t+1}^{1},\ldots,a_{t+1}^{N}\mid\phi^{-}\right)|_{a_{t+1}^{n}=\pi\left(o_{t+1}\right),n=1,\ldots,N}$ (5)
@@
The shared reward makes the agents learn policy cooperatively. In the execution phase, we only use the policy $\pi$ whose inputs are $o_{t}^{1},\ldots,o_{t}^{N}$, respectively, and the outputs are force ($\boldsymbol{F}$) applied to the agent. According to the physical dynamics of agents described in Section 4.1, we can get the actions $\boldsymbol{a}_{t}=\left(a_{t}^{1},\ldots,a_{t}^{N}\right)$. Note that, for agent $n$, the input of $\pi$ is only the observation $o_{t}^{n}$ received by agent $n$ at time step $t$ in the execution phase.
5 Experiments
In this section, we verified our method from three aspects. In Section 5.1, we verified the target assignment and path planning performances of our method. In all the experiments, the locations of agents and task points are generated randomly. In Section 5.2, we verified the learned cooperation ability. In Section 5.3, we verified the time efficiency.
5.1 Target Assignment and Path Planning
In this subsection, we verified our method in various of intelligent warehouse settings. We set five different level scenarios to verify the performances: (1) two agents - two tasks (2) two agents - four tasks (3) five agents - five tasks (4) five agents - ten tasks (5) five agents - twenty tasks.
Figure 5 shows the training average return curves, in all the different level scenarios, the average returns increase monotonously. This verifies the stability of our method.
Figure 6 to Figure 10 show the performance target assignment and pathfinding in five different level scenarios. Experimental results show that although the difficulty of these five different level scenarios increases gradually, our method performs well in all the tasks.
(a) two agents - two tasks
(b) two agents - four tasks
(c) five agents - five tasks
(d) five agents - ten tasks
(e) five agents - twenty tasks
@@ 5.1 Target Assignment and Path Planning
For easy tasks (two agents - two tasks) in Figure 6 (a) (b) (c) (d), we can see that the target assignment and pathfinding were addressed well. For target assignment, results show that the task assignment is addressed very reasonably because our method assigns the task to the agent that is close to the task. For pathfinding, it is shown that the planned paths are almost shortest.
For difficult tasks (five agents - twenty tasks) in Figure 10 (a) (b) (c) (d), we can see that the target assignment and pathfinding were also addressed well. For target assignment, it can be seen that each agent is reasonably assigned to several nearby tasks. For pathfinding, it is shown that the planned paths are usually the shortest. The same superior performances can be seen in other different level scenarios, such as two agents - four tasks in Figure 7, five agents - five tasks in Figure 8, and five agents - ten tasks in Figure 9. This demonstrated the efficiency of our method.
@@
Figure 11: The cooperation of agents
Table 1: Time comparisons between our method and traditional methods
@@
5.2 Cooperation Ability
We also designed a conflict scenario to verify the learned cooperation ability of agents. As shown in Figure 11, the big purple circle represents agent-1, its task is denoted by the small purple circle. The big gray circle represents agent-2, its task is denoted by the small black circle. We deliberately block other roads to create a conflict environment. As shown in Figure 11, there is bound to be a conflict between agent-1 and agent-2 during the navigation. The trajectories generated by our method are shown in Figure 11, the red curve is the trajectory of agent-1, and the cyan curve is the trajectory of agent-2. Results show that both agent-1 and agent-2 had learned to avoid each other at the point of conflict, and then navigated to their tasks. This verified the cooperation ability of our method.
5.3 Time Efficiency
@@
Table 1 shows the time comparisons between traditional methods [35, 36] and our method. Table 1 shows that our method simultaneously addresses target assignment (TA) and path planning (PF) problems, however, the traditional method addresses target assignment first and then performs path planning. Thus, the consumed time in our method is TAPF time, however, the consumed time in the traditional method is TA+PF. In Table 1, ES means each step. The reasons that we compare the consumed time of each step is: (1) Our method handles continuous action space problems, however, the traditional method handles discrete action space problems. (2) The consumed time of path finding can be influenced by many factors, such as the resolution of the grid map and the distance taken at each step in the grid map. (3) previous literature [35, 36] did not consider the physical dynamics of agents. However, the physical dynamics of agents have been considered in this study. Thus, for a fair comparison, we only compare the consumed time of each step, which means the consumed time from an agent obtains an observation to output the decision action. For the traditional method, the consumed time of target assignment is computed by the method in [35], and the consumed time of each step in path planning is computed by the method in [36].
@@
Table 1 shows that in easy tasks such two agents - two tasks and two agents - four tasks, the consumed time of target assignment in the traditional method can be acceptable. However, as the difficulty of tasks increased, the time of solving target assignments in the traditional method increased rapidly, especially in the five agents - twenty tasks scenario. However, our method provides policy time-efficiently in all different level tasks (from two agents - two tasks to five agents - twenty tasks). This verified the time efficiency ability of our method. As we all know, time efficiency is an important factor in real-world TAPF problems. Therefore, the experiment results demonstrated the promising engineering practical value of our method. Note that Figure 6 to Figure 10 had shown the performances of target assignment and path finding of our method in five different level scenarios. For target assignment, results show that the task assignment is addressed reasonably because our method assigns the task to the agent that is close to the task. For pathfinding, results show that the planned paths are almost shortest.
6 Conclusion and future work
@@
In this study, we propose a method to simultaneously address the TAPF problem in the intelligent warehouse from a perspective of cooperative multi-agent deep RL. First, we model the TAPF problem as a cooperative multi-agent deep RL problem. Then, we simultaneously address target assignment and pathfinding by cooperative multi-agent deep RL algorithm. Moreover, previous literature rarely considers the physical dynamics of agents. In this study, the physical dynamics of agents have been considered. Experimental results show that our method performs well in various task settings, which means the target assignment is solved reasonably and the planned path is almost the shortest. Furthermore, our method is more time-efficient than baselines. For future work, we will apply our method to real-world target assignments and pathfinding problems.
Appendix
Network Architecture: The actor network contains four fully connected layers. As described in Section 4.1, the input dimension of the first layer is equal to the observation dimension. The hidden dimension is 128. The first three layers are followed by ReLU activation function. The output dimension of the fourth layer is the action dimension. The critic network contains four fully connected layers. The input dimension of the first layer is equal to the observation dimension plus the action dimension. The hidden dimension is 128. The first three layers are followed by ReLU activation function. The output of the fourth layer is the state action value. Networks are trained by Adam optimizer.
[…]

### v3

@@ IV-A Modeling TAPF as a MARL Problem
The physical dynamics of agents: Before describing the action space, we first describe the dynamics of agents. Note that our work is the first to consider the dynamics of agents in TAPF for the intelligent warehouse. As shown in Fig. 2 (a), the "No dynamics" schematic diagram represents previous literature [29, 35] methods that only focus on the next position of an agent but do not consider velocities and accelerations of agents. Further, the dynamics of agents in previous literature are solved by using traditional methods [36]. On the contrary, the "Dynamics" schematic diagram in Fig. 2 (b) represents our method that considers the dynamics of agents. We can calculate agents’ velocities and accelerations.
The output of our policy network is the magnitude and direction of four cardinal directions force ($\vec{F}_{x},\vec{F}_{-x},\vec{F}_{y},\vec{F}_{-y}$) applied to the agent. According to Newton’s second law of motion, $\vec{F}_{direction}=m*\vec{a}_{direction}$ (where $m$ is the mass of the agent), we can get the acceleration $\vec{a}_{direction}=\frac{\vec{F}_{direction}}{m}$. According to basic physics knowledge, we can get the velocity of the agent $\vec{v}_{t(direction)}=\vec{v}_{0(direction)}+\vec{a}_{direction}*\Delta t$, where $v_{0}$ is the initial velocity, and $\Delta t$ denotes a time interval.
Action space: In this study, the action space of an agent is continuous, representing the movements. As shown in Fig. 3, an agent obtains a velocity between 0.0m/s and 1.0m/s in each of the four cardinal directions [move_left ($\vec{v}_{-x}$), move_right ($\vec{v}_{x}$), move_down ($\vec{v}_{-y}$), move_up ($\vec{v}_{y}$)], the final action ($\vec{v}$) is calculated by the vector sum of the four velocities.
Reward function: The reward function is defined as:
@@
Figure 4: Distansce matrix of tasks and agents: $\|\bm{P_{tasks}-P_{agents}}\|$. The element $dis(a\ i,t\ i)$ represents the distance between $agent\ i$ and $task\ i$.
@@
    wherendenotes the number of elements less than 0.05 in
@@
  • reward_distance_tasks_to_agents
    = -$\|\bm{P_{tasks}-P_{agents}}\|_{min}$
  • reward_collision_agents_to_obstacles = -$2*n$,
    wherendenotes the number of elements less than R in
@@
  • reward_collision_agents_to_agents = -$2*n$,
    wherendenotes the number of elements less than 2R in
[…]
For reward_success, $\|\bm{P_{tasks}-P_{agents}}\|$ represents the distance matrix of all tasks to all agents. Fig. 4 shows a detailed distance matrix. If the distance of $agent\ i$ to $task\ i$ is less than $0.05m$, we give a $+100$ positive reward to force agents to navigate to the tasks. If there are $n$ elements in $\|\bm{P_{tasks}-P_{agents}}\|$ less than $0.05m$, we set reward_success = +$100*n$. Reward_distance_tasks_to_agents is created to provide a dense reward to accelerate the training speed. In $\|\bm{P_{tasks}-P_{agents}}\|_{min}$, the subscript $min$ means that we sum the minimum distance of each task to every agent.
Reward_collision_agents_to_obstacles aims to punish agents from collisions with obstacles. The definition of $\|\bm{P_{obstacles}-P_{agents}}\|$ is similar with $\|\bm{P_{tasks}-P_{agents}}\|$, $\|\bm{P_{obstacles}-P_{agents}}\|$ denotes the distance matrix of obstacles and tasks. If the distance of an agent to an obstacle is less than $R$, a negative reward $-2$ is given. If there are $n$ elements in $\|\bm{P_{obstacles}-P_{agents}}\|$ less than $R$, we set reward_collision_agents_to_obstacles = -$2*n$.
Reward_collision_agents_to_agents aims to punish agents from collision to agents. The definition of $\|\bm{P_{agents}-P_{agents}}\|$ is similar with $\|\bm{P_{tasks}-P_{agents}}\|$, $\|\bm{P_{agents}-P_{agents}}\|$ denotes the distance matrix of agents and agents. If the distance of an agent to an agent is less than $2R$, we give a negative reward $-2$. If there are $n$ elements in $\|\bm{P_{agents}-P_{agents}}\|$ less than $2R$, we set reward_collision_agents_to_agents = -$2*n$.
IV-B Solving TAPF Problem via Cooperative MARL
In this study, we introduce MADDPG [33] to solve the TAPF problem. Considering that agents are homogeneous, they can share the same policy network. This makes learning policy more efficient. Let $N$ denote the number of agents, and $\pi$ denote the policy parameterized by $\theta$. The centralized critic $Q$ is parameterized by $\phi$. The parameters $\theta$ for policy $\pi$ can be updated iteratively by the associated critic $Q$:
$\nabla_{\theta}J\left(\theta\right)$ $=\mathbb{E}_{\left(\bm{x_{t}},\bm{a_{t}}\right)\sim\mathcal{D}}\Big{[}\nabla_{\theta}\pi\left(o_{t}^{n}\mid\theta\right)$ (3)
$\quad\nabla_{a_{t}}Q\left(\bm{x_{t}},a_{t}^{1},\ldots,a_{t}^{N}\mid\phi\right)|_{a_{t}^{n}=\pi\left(o_{t}^{n}\mid\theta\right)}\Big{]}$
where $\bm{x}_{t}$ and $\bm{a}_{t}$ are the concatenation of all agents’ observation and action at time step $t$, $\bm{x}_{t}=\left(o_{t}^{1},\ldots,o_{t}^{N}\right)$ and $\bm{a}_{t}=\left(a_{t}^{1},\ldots,a_{t}^{N}\right)$. $o_{t}^{n}$ is the observation received by agent $n$ at time step $t$. $\mathcal{D}$ denotes the replay buffer containing tuples ($\bm{x}_{t},\bm{a}_{t},r_{t}\cdots$).
@@ IV-B Solving TAPF Problem via Cooperative MARL
$L\left(\phi\right)=\mathbb{E}_{\left(\bm{x}_{t},\bm{a}_{t},{r}_{t},\bm{x}_{t+1}\right)\sim\mathcal{D}}\left[\left(Q\left(\bm{x}_{t},\bm{a}_{t}\mid\phi\right)-y_{t}\right)^{2}\right]$ (4)
where ${r}_{t}$ is the concatenation of rewards (Eq (2)) received by all agents at time step $t$. The target critic value $y_{t}^{n}$ is defined as:
$y_{t}=r_{t}+\gamma Q^{-}\left(\bm{x}_{t+1},a_{t+1}^{1},\ldots,a_{t+1}^{N}\mid\phi^{-}\right)|_{a_{t+1}^{n}=\pi\left(o_{t+1}\right),n=1,\ldots,N}$ (5)
@@
The shared reward makes the agents learn policy cooperatively. In the execution phase, we only use the policy $\pi$ whose inputs are $o_{t}^{1},\ldots,o_{t}^{N}$, respectively, and the outputs are force ($\bm{\vec{F}}$) applied to the agent. According to the physical dynamics of agents described in Section IV-A, we can get the actions $\bm{a}_{t}=\left(a_{t}^{1},\ldots,a_{t}^{N}\right)$. Note that, for agent $n$, the input of $\pi$ is only the observation $o_{t}^{n}$ received by agent $n$ at time step $t$ in the execution phase.
@@
V Experiments
In this section, we verified the proposed method from three aspects. In Section V-A, we verified the proposed method’s target assignment and path planning performances. In all the experiments, the locations of agents and task points are generated randomly. Section V-B verified the learned cooperation ability. Section V-C verified the time efficiency.
@@ V Experiments
V-A Target Assignment and Path Planning
This subsection verified the proposed method in various intelligent warehouse settings. We set five different level scenarios to verify the performances: (1) two agents - two tasks (2) two agents - four tasks (3) five agents - five tasks (4) five agents - ten tasks (5) five agents - twenty tasks.
Fig. 5 shows the training average return curves. In all the different level scenarios, the average returns increase monotonously. This verifies the stability of the proposed method. Fig. 6 to Fig. 10 show the performance target assignment and pathfinding in five different level scenarios. Experimental results show that although the difficulty of these five different level scenarios increases gradually, the proposed method performs well in all the tasks.
@@ V-A Target Assignment and Path Planning
For easy tasks (two agents - two tasks) in Fig. 6 (a) (b) (c) (d), we can see that the target assignment and pathfinding were addressed well. For target assignment, results show that the task assignment is addressed very reasonably because the proposed method assigns the task to the agent that is close to the task. For pathfinding, it is shown that the planned paths are almost shortest.
@@
For difficult tasks (five agents - twenty tasks) in Fig. 10 (a) (b) (c) (d), we can see that the target assignment and pathfinding were also addressed well. For target assignment, it can be seen that each agent is reasonably assigned to several nearby tasks. For pathfinding, it is shown that the planned paths are usually the shortest. The same superior performances can be seen in other different level scenarios, such as two agents - four tasks in Fig. 7, five agents - five tasks in Fig. 8, and five agents - ten tasks in Fig. 9. This demonstrated the efficiency of the proposed method.
@@
V-B Cooperation Ability
We also designed a conflict scenario to verify the agents’ learned cooperation ability. As shown in Fig. 11, the big purple circle represents agent-1, its task is denoted by the small purple circle. The big gray circle represents agent-2, and the small black circle denotes its task. We deliberately block other roads to create a conflict environment. As shown in Fig. 11, there is bound to be a conflict between agent-1 and agent-2 during the navigation. The trajectories generated by the proposed method are shown in Fig. 11. The red curve is the trajectory of agent-1, and the cyan curve is the trajectory of agent-2. Results show that both agent-1 and agent-2 learned to avoid each other at the point of conflict and then navigated to their tasks. This verified the cooperation ability of the proposed method.
TABLE I: Time comparisons between our method and traditional methods
@@ V-B Cooperation Ability
Figure 11: The cooperation of agents
V-C Time Efficiency
@@
Table I shows the time comparisons between traditional methods [37, 38] and the proposed method. Table I shows that the proposed method simultaneously addresses target assignment (TA) and path planning (PF) problems. However, the traditional method addresses target assignment first and then performs path planning. Thus, the time consumed by the proposed method is TAPF time. However, the time consumed by the traditional method is TA+PF. In Table I, ES means each step. The reasons that we compare the consumed time of each step are: (1) Our method handles continuous action space problems. However, the traditional method handles discrete action space problems. (2) The consumed time of path finding can be influenced by many factors, such as the resolution of the grid map and the distance taken at each step in the grid map. (3) Previous literature [37, 38] did not consider the physical dynamics of agents. However, the physical dynamics of agents have been considered in this study. Thus, for a fair comparison, we only compare the consumed time of each step, which means the consumed time from an agent obtains an observation to output the decision action. For the traditional method, the consumed time of target assignment is computed by the method in [37], and the consumed time of each step in path planning is computed by the method in [38].
@@
Table I shows that in easy tasks such two agents - two tasks and two agents - four tasks, the consumed time of target assignment in the traditional method can be acceptable. However, as the difficulty of tasks increased, the time of solving target assignments in the traditional method increased rapidly, especially in the five agents - twenty tasks scenario. However, the proposed method provides policy time-efficiently in all different level tasks (from two agents - two tasks to five agents - twenty tasks). This verified the time efficiency ability of the proposed method. As we all know, time efficiency is an important factor in real-world TAPF problems. Therefore, the experiment results demonstrated the promising engineering practical value of the proposed method. Note that Fig. 6 to Fig. 10 had shown the performances of target assignment and path finding of the proposed method in five different level scenarios. For target assignment, results show that the task assignment is addressed reasonably because the proposed method assigns the task to the agent that is close to the task. For pathfinding, results show that the planned paths are almost shortest.
VI Conclusion and future work
@@
This paper proposed a method to simultaneously solve the TAPF problem in the intelligent warehouse from a cooperative multi-agent deep RL perspective. First, we modeled the TAPF problem as a cooperative multi-agent deep RL problem. Then, we simultaneously solved task assignment and path finding using a cooperative multi-agent deep RL algorithm. Moreover, previous literature rarely considers the physical dynamics of agents. In this study, the physical dynamics of agents have been considered. Experimental results showed that the proposed method performed well in various task settings, which meant the target assignment was solved reasonably, and the planned path was almost the shortest. Furthermore, the proposed method is more time-efficient than baselines. For future work, we will apply the proposed method to real-world TAPF problems.